# The influence of Parker spiral on the reflection-driven turbulence

### v1

@@ 2.4 Reflection-driven Turbulence Phenomenology
The RDT framework rests on following assumptions that demand careful scrutiny. First, the nonlinear damping rate provided by the turbulence is ${z}^{\mp}/\ell_{\perp}$, implying strong turbulence ($\chi_{\rm A}\sim 1$) for validity (but see related weak turbulence version in Chandran2019), which leads to the same scaling. Second, the phenomenology assumes that a single dominant scale $\ell_{\perp}$ governs both $\tilde{z}^{+}$ and $\tilde{z}^{-}$; this $\ell_{\perp}$ cancels in the heating rate $\mathcal{Q}$, but in principle could be different for $\tilde{z}^{+}$ and $\tilde{z}^{-}$ (indeed, this was seen by Meyrand2025). Third, anomalous coherence, with $\tilde{z}^{-}$ carried along by $\tilde{z}^{+}$, is required to preserve the $\tilde{z}^{-}$ balance and to justify neglecting the ($v_{\rm A}\cdot\nabla$) term in our first assumption. Fourth, $\tilde{z}^{-}$ must adjust to reflection-nonlinear equilibrium faster than background changes; this requirement reduces to $\chi_{\rm exp}\gtrsim 1$. Fifth, reflection in the $\tilde{z}^{+}$ equation must be negligible ($\mathcal{R}^{+}\tilde{z}^{-}\ll\mathcal{N}^{+}\tilde{z}^{+}$), which likewise reduces to the same $\chi_{\rm exp}\gtrsim 1$ condition.
@@
Thus the phenomenology is self-consistent only for sufficiently large amplitude $\tilde{z}^{+}$ such that $\chi_{\rm exp}\gg 1$, which implies strong imbalance $\tilde{z}^{-}\ll\tilde{z}^{+}$. We therefore expect the system to transition to the balanced regime once the outward Elsässer energy decays to the point that $\chi_{\rm exp}\sim 1$. Conversely, when $\chi_{\rm exp}<1$, the ordering breaks down, reflection becomes dominant and the dynamics from (16) likely revert toward reflection-dominated (essentially linear) behavior. In this regime with a radial $\bar{\bm{B}}$, the system evolves into a magnetically dominated, balanced state in which $\tilde{\bm{z}}^{+}\simeq-\tilde{\bm{z}}^{-}$, and $\tilde{E}$ starts growing in time (see Eq. 10). Nonlinear dissipation is strongly suppressed, so the turbulent heating effectively shuts off (Meyrand2025).
@@ 2.4.1 The role of Parker spiral
Figure 1: Schematic of an expanding plasma parcel in the solar wind. (Top) Geometry of the expanding box with axes aligned to radial ($x$), azimuthal ($y$), and normal ($z$) directions. The solar wind initially flows radially outward, while the box expands transversely. The mean magnetic field $\bar{\bm{B}}$ (green) is radial near the Sun but rotates into a Parker spiral with angle $\Phi$ as $a(t)$ increases, so that $B_{y}/B_{x}\sim a(t)$. Wavevectors $\bm{k}$ are shown at angle $\vartheta$ to $\bar{\bm{B}}$, with azimuthal ($k^{(Y)}$) and normal ($k^{(Z)}$) components indicated. (Bottom) Comparison of eddy evolution under (a) purely radial and (b) Parker-spiral expansion. In the radial case, the perpendicular scale $\ell_{\perp}$ grows uniformly with expansion. In the PS case, rotation of the mean field changes $\ell_{\perp}$ creating 3-D anisotropy of the eddies.
@@
Figure 2: Evolution of key parameters for waves under solar-wind expansion, plotted versus expansion factor $a$. Thick black: radial field ($\Phi_{0}=0^{\circ}$). Colored: Parker spiral cases ($\Phi_{0}=2^{\circ}\text{--}20^{\circ}$, darker colors correspond to larger $\Phi_{0}$). (a) Wave amplitude $z^{+}/v_{\rm A}$ stays roughly constant for the purely radial case, and in the Parker spiral (PS) case remains constant initially but then decays as $\propto a^{-1}$ once the azimuthal component becomes significant and $v_{\rm A}\propto a^{0}$. (b) Obliquity $\sin\vartheta(a)$: the angle between $\bm{k}$ and $\bar{\bm{B}}$ initially decreases and then increases again as the mean field rotates azimuthally, producing a clear inflection with a PS. (c) Expansion–cascade parameter $\chi_{\rm exp}=(k_{\perp}\,z^{+}_{\perp})/(\dot{a}/a)$ for the out-of-plane case ($\varphi=\pi/2$): turbulence is sustained while $\chi_{\rm exp}\gtrsim 1$; $\chi_{\rm exp}$ initially decays as $\propto a^{-1}$ for both radial and Parker spiral cases, but for PS it starts increasing and eventually flattens as $a^{0}$ once the azimuthal component becomes dominant. (d) $\chi_{\rm exp}$ for in-plane wave vectors: shown for $\varphi=\pi$ (solid) and $\varphi=0$ (dashed); in both cases, the wave vector lies in the xxyy plane. For $\varphi=0$, the wave passes through purely parallel propagation.
@@
In the Parker spiral case, $\sin\vartheta$ initially decreases as in the radial case, but as $a$ approaches $a_{\min}=\sqrt{\cot\theta_{p0}\,\cot\Phi_{0}}$, its evolution deviates: $\sin\vartheta$ reaches a minimum and then increases again as $\overline{\bm{B}}$ rotates toward the perpendicular direction (see Fig. 2b and also Squire2022). While the detailed dynamics are complex and nonlinear, we assume the overall RDT framework still applies such that $z^{+}\propto a^{-1}$, but the evolving geometry changes the scaling of $k_{\perp}$. Initially $k_{\perp}$ decays with expansion as in the radial case, but as the mean field rotates relative to the eddies, making wave vectors oblique and driving $k_{\perp}$ back up, thereby increasing the non-linear turnover rate. As a result, instead of continuously decaying, $\chi_{\rm exp}$ levels off staying roughly constant ($\propto a^{0}$) or even growing at intermediate times, delaying indefinitely the turbulence shutoff that occurs at $\chi_{\rm exp}<1$ and maintaining the high imbalance (see Fig. 2).
@@ 3.1 Expanding box model numerical
To solve the EBM Eqs. (1)–(3), we use the finite-volume astrophysical code Athena++ (Stone2020). We employ the Harten-Lax-van Leer Discontinuities (HLLD) Riemann solver (Mignone2007), modified to include expansion effects. To simplify the expansion terms and enhance numerical stability, we transform variables as follows (Johnston2022):
@@ 3.2 Simulations setup and initial conditions
Numerical details of all simulations are given in Table 1. Each simulation is performed in an initially elongated box that co-moves with the solar wind under the expanding-box framework. We set the initial box dimensions to $L_{x0}=1$ and $L_{y0}=L_{z0}=0.2$, starting from an anisotropic domain to capture transverse stretching as plasma advects outward. We will use expansion factor $a$, rather than time in discussing all results. As the system expands, the box elongates in the transverse directions ($L_{y}$, $L_{z}$ $\propto a$); by $a=50$, the domain becomes strongly oblate, resembling a pancake-like structure. At $a=1$, we initialize the background magnetic field $\bar{\bm{B}}$} in the $xy$ plane with $|{\bar{\bm{B}}}|=1$ and $B_{y,0}<0$, defining the initial Parker‐spiral angle $\Phi_{0}$.
@@
which measures how the components of $\bm{B}$ are correlated to keep $|\bm{B}|$ constant. $C_{B}^{2}=0$ for perfectly spherically polarized fluctuations. The initial condition is chosen to mimic a strongly $z^{+}$-dominated, nearly spherically polarized Alfvénic wind propagating outwards from near the Alfvén point: $\bm{z}^{+}(t=0)$ carries the fluctuation energy while $\bm{z}^{-}(t=0)=0$. Radial expansion and large-scale inhomogeneity produce partial reflection (creation of $\bm{z}^{-}$) as the packet propagates outward, thereby initiating reflection-driven turbulence. We initialize the fluctuations as linearly polarized Alfvén waves $(\bm{k}\times{{\bar{\bm{B}}}})$ (Squire2020; Johnston2022), and the modes rapidly (within $\sim 0.5$ Alfvén time) lose the variation in $|\bm{B}|$ caused by their initial linear polarization, reducing $C_{B}^{2}$ to quickly become spherically polarized.
The initial set of Alfvén waves are generated from a sum of randomly phased waves with amplitude determined by Gaussian energy spectrum,
@@
Here $k_{\parallel,0}=\kappa_{\parallel}({2\pi}/{L_{x}}),\;k_{\perp,0}=\kappa_{\perp}({2\pi}/{L_{\perp}}),\;\text{and}\;k_{w}=({12}/{L_{\perp}})$, which set the scale of the center of Gaussian peak and the width of the peak respectively. $k_{\parallel,0}$ and $k_{\perp,0}$ lie respectively parallel and perpendicular to $\bar{\bm{B}}$, maintaining anisotropy even with an initial tilted spiral field.
To choose simulation parameters and quantify regimes, we use the timescales and their ratios, as discussed in section˜2. The expansion time is $\tau_{\rm exp}^{-1}=\dot{a}/a$, the linear timescale is $\tau_{\rm A}^{-1}=k_{\parallel}v_{\rm A}$, and the outer-scale nonlinear time is $\tau_{\rm nl}^{-1}=k_{\perp}z^{+}$. From these rates we form the dimensionless control parameters: (i) $\chi_{\rm exp,0}=(k_{\perp,0}\;z^{+}_{\rm rms0})/(\dot{a}/a)$; (ii) $\chi_{\rm A}\doteq(k_{\perp}z^{+})/(k_{\parallel}v_{\rm A})$, which compares strength of nonlinear interactions to linear Alfvénic propagation and thus quantifies how strongly nonlinearity competes with Alfvénic decorrelation of $z^{-}$ fluctuations (GS95); and (iii) the ratio of box-scale Alfvén time to the expansion time $\Delta_{0}\doteq\chi_{\rm exp0}/\chi_{\rm A0}=v_{\rm A,0}/\dot{a}L_{x}$ which is constant. For the parameter set used in this paper we fix $\Delta_{0}=2$ (via $\dot{a}=0.5$, $L_{\parallel}=1$ and $v_{\rm A}=1$) and vary $\chi_{\rm exp,0}$ roughly in the range $\approx 11\!-\!30$. We explore two initial fluctuation amplitudes defined by $A\equiv{z^{+}_{\rm rms0}}/{v_{\rm A,0}}=1.0\;\text{and}\;0.5,$ where $z^{+}_{\rm rms0}$ is the initial root-mean-square fluctuation amplitude (see Table 1 for exact values). Changing $A$ alters the initial nonlinear time (and therefore $\chi_{\rm exp,0}$ and $\chi_{\rm A,0}$), so we fix $\chi_{\rm exp,0}$ to isolate pure amplitude effects. Physically, larger $A$ produces larger fluctuations with larger $\delta B_{\parallel}$ to maintain their spherical polarization, whereas smaller $A$ (initial conditions) remain closer to the reduced MHD limit, which is derived assuming $z^{+}/v_{\rm A}\ll 1$.
@@ 4.1 Turbulence Evolution and Emergence of the Parker Spiral
The morphology of turbulence depends sensitively on the background magnetic geometry. The top panels of Fig. 3a depict two-dimensional snapshots ($y$-$z$ mid plane slice) of the perpendicular normalized Elsässer fields $|\bm{z}_{\perp}^{\pm}|$ for the radial A05-$\Phi_{0}\!=\!0^{\circ}$ run (here $\bm{z}_{\perp}^{\pm}=(\bm{I}-\hat{\bm{b}}\hat{\bm{b}})\cdot\bm{z}^{\pm}$, which is approximately the Alfvénic part). In the radial case, turbulence evolves in two distinct phases. During the early, imbalanced stage when $\chi_{\rm exp}>1$, outward $z^{+}$-fluctuations are strong and develop fine perpendicular structure, while reflected $z^{-}$ fluctuations appear as more intermittent, field-aligned filaments. As expansion proceeds and $\chi_{\rm exp}$ approaches unity, reflected normalized wave-action energy grows and $\tilde{E}^{-}/\tilde{E}^{+}$ approaches unity, with both fields acquiring increasingly similar circular vortex-like structures. In this balanced phase, the nonlinear transfer rate decreases relative to expansion and the cascade stalls, leaving coherent vortices to dominate the dynamics and shutting off heating. Throughout both stages, expansion stretches eddies across the mean field, generating large-aspect-ratio quasi 2D eddies towards the simulation’s end. This behavior qualitatively matches the reduced-MHD results of Meyrand2025, who reported 2D “Alfvén vortices”, similar to those we see, in the late stage $\chi_{\rm exp}\sim 1$ magnetically dominated regime.
@@
The Parker spiral simulations evolve differently. In A05-$\Phi_{0}\!=\!1.5^{\circ}$ (lower two panels of Fig. 3), turbulence does not relax toward a quasi-two-dimensional state as in the radial case. Instead, nonlinear activity persists to larger expansion factors because (as we will show below) $\chi_{\rm exp}$ remins larger. The fields develop persistent, ribbon-like structures with sharper transverse gradients and smaller $\ell_{\perp}$ than in radial runs (this will be quantified in section˜4.3). Physically, this follows because the continuous rotation of the mean field during expansion maintains a persistent misalignment between the perpendicular and expansion directions; as illustrated in Fig. 1 and discussed in section˜2.4.1, that misalignment prevents eddies from maintaining cylindrical symmetry about the local field and thereby inhibits the formation of quasi-2D vortices. Consequently, we observe a persistent, irregular 3D anisotropy in PS runs and persistent turbulent heating.
@@
Figure 4 shows similar features via a three-dimensional depiction of the Elsässer fields from the high-resolution (HR-$\Phi_{0}\!=\!0^{\circ}$ and HR-$\Phi_{0}\!=\!4^{\circ}$) simulations. These volumetric views support the conclusion of Fig. 3, and also show the evolution of radial structure. In HR-$\Phi_{0}\!=\!0^{\circ}$, turbulence condenses into slowly developing Alfvén vortices, though these are not so obvious here because of the earlier time. In HR-$\Phi_{0}\!=\!4^{\circ}$, structures stay at smaller scale and are possibly more intermittent, developing 3D outer-scale anisotropy. The resulting sustained small $\ell_{\perp}$ inhibits the cascade from stalling. We also see qualitative agreement with prior studies linking the PS geometry to 3D anisotropy (Verdini2019).

### v2

@@ 2.4 Reflection-driven Turbulence Phenomenology
The RDT framework rests on following assumptions that demand careful scrutiny. First, the nonlinear damping rate provided by the turbulence is ${z}^{\mp}/\ell_{\perp}$, implying strong turbulence ($\chi_{\rm A}\sim 1$) for validity (but see related weak turbulence version in Chandran and Perez 2019), which leads to the same scaling. Second, the phenomenology assumes that a single dominant scale $\ell_{\perp}$ governs both $\tilde{z}^{+}$ and $\tilde{z}^{-}$; this $\ell_{\perp}$ cancels in the heating rate $\mathcal{Q}$, but in principle could be different for $\tilde{z}^{+}$ and $\tilde{z}^{-}$ (indeed, this was seen by Meyrand et al. 2025). Third, anomalous coherence, with $\tilde{z}^{-}$ carried along by $\tilde{z}^{+}$, is required to preserve the $\tilde{z}^{-}$ balance and to justify neglecting the ($v_{\rm A}\cdot\nabla$) term in our first assumption. Fourth, $\tilde{z}^{-}$ must adjust to reflection-nonlinear equilibrium faster than background changes; this requirement reduces to $\chi_{\rm exp}\gtrsim 1$. Fifth, reflection in the $\tilde{z}^{+}$ equation must be negligible ($\mathcal{R}^{+}\tilde{z}^{-}\ll\mathcal{N}^{+}\tilde{z}^{+}$), which likewise reduces to the same $\chi_{\rm exp}\gtrsim 1$ condition.
@@
Thus the phenomenology is self-consistent only for sufficiently large amplitude $\tilde{z}^{+}$ such that $\chi_{\rm exp}\gg 1$, which implies strong imbalance $\tilde{z}^{-}\ll\tilde{z}^{+}$. We therefore expect the system to transition to the balanced regime once the outward Elsässer energy decays to the point that $\chi_{\rm exp}\sim 1$. Conversely, when $\chi_{\rm exp}<1$, the ordering breaks down, reflection becomes dominant and the dynamics from (16) likely revert toward reflection-dominated (essentially linear) behavior. In this regime with a radial $\bar{\bm{B}}$, the system evolves into a magnetically dominated, balanced state in which $\tilde{\bm{z}}^{+}\simeq-\tilde{\bm{z}}^{-}$, and $\tilde{E}$ starts growing in time (see Eq. 10). Nonlinear dissipation is strongly suppressed, so the turbulent heating effectively shuts off (Meyrand et al., 2025).
@@ 2.4.1 The role of Parker spiral
Figure 1: Schematic of an expanding plasma parcel in the solar wind. (Top) Geometry of the expanding box with axes aligned to radial ($x$), azimuthal ($y$), and normal ($z$) directions. The solar wind initially flows radially outward, while the box expands transversely. The mean magnetic field $\bar{\bm{B}}$ (green) is radial near the Sun but rotates into a Parker spiral with angle $\Phi$ as $a(t)$ increases, so that $B_{y}/B_{x}\sim a(t)$. Wave-vectors $\bm{k}$ are shown at angle $\vartheta$ to $\bar{\bm{B}}$, with azimuthal ($k^{(Y)}$) and normal ($k^{(Z)}$) components indicated. (Bottom) Comparison of eddy evolution under (a) purely radial and (b) Parker-spiral expansion. In the radial case, the perpendicular scale $\ell_{\perp}$ grows uniformly with expansion. In the PS case, rotation of the mean field changes $\ell_{\perp}$ creating 3-D anisotropy of the eddies.
@@
Figure 2: Evolution of key parameters for waves under solar-wind expansion, plotted versus expansion factor $a$. Thick black: radial field ($\Phi_{0}=0^{\circ}$). Colored: Parker spiral cases ($\Phi_{0}=2^{\circ}\text{--}20^{\circ}$, darker colors correspond to smaller $\Phi_{0}$). (a) Wave amplitude $z^{+}/v_{\rm A}$ stays roughly constant for the purely radial case, and in the Parker spiral (PS) case remains constant initially but then decays as $\propto a^{-1}$ once the azimuthal component becomes significant and $v_{\rm A}\propto a^{0}$. (b) Obliquity $\sin\vartheta(a)$: the angle between $\bm{k}$ and $\bar{\bm{B}}$ initially decreases and then increases again as the mean field rotates azimuthally, producing a clear inflection with a PS. (c) Expansion–cascade parameter $\chi_{\rm exp}=(k_{\perp}\,z^{+}_{\perp})/(\dot{a}/a)$ for the out-of-plane case ($\varphi=\pi/2$): turbulence is sustained while $\chi_{\rm exp}\gtrsim 1$; $\chi_{\rm exp}$ initially decays as $\propto a^{-1}$ for both radial and Parker spiral cases, but for PS it starts increasing and eventually flattens as $a^{0}$ once the azimuthal component becomes dominant. (d) $\chi_{\rm exp}$ for in-plane wave vectors: shown for $\varphi=\pi$ (solid) and $\varphi=0$ (dashed); in both cases, the wave vector lies in the xxyy plane. For $\varphi=0$, the wave passes through purely parallel propagation.
@@
In the Parker spiral case, $\sin\vartheta$ initially decreases as in the radial case, but as $a$ approaches $a_{\min}=\sqrt{\cot\theta_{p0}\,\cot\Phi_{0}}$, its evolution deviates: $\sin\vartheta$ reaches a minimum and then increases again as $\overline{\bm{B}}$ rotates toward the perpendicular direction (see Fig. 2b and also Squire et al. 2022). While the detailed dynamics are complex and nonlinear, we assume the overall RDT framework still applies such that $z^{+}\propto a^{-1}$, but the evolving geometry changes the scaling of $k_{\perp}$. Initially $k_{\perp}$ decays with expansion as in the radial case, but as the mean field rotates relative to the eddies, making wave vectors oblique and driving $k_{\perp}$ back up, thereby increasing the non-linear turnover rate. As a result, instead of continuously decaying, $\chi_{\rm exp}$ levels off staying roughly constant ($\propto a^{0}$) or even growing at intermediate times, delaying indefinitely the turbulence shutoff that occurs at $\chi_{\rm exp}<1$ and maintaining the high imbalance (see Fig. 2).
@@ 3.1 Expanding box model numerical
To solve the EBM Eqs. (1)–(3), we use the finite-volume astrophysical code Athena++ (Stone et al., 2020). We employ the Harten-Lax-van Leer Discontinuities (HLLD) Riemann solver (Mignone, 2007), modified to include expansion effects. To simplify the expansion terms and enhance numerical stability, we transform variables as follows (Johnston et al., 2022):
@@ 3.2 Simulations setup and initial conditions
Numerical details of all simulations are given in Table 1. All simulations employ periodic boundary conditions in $x,y,z$, as is standard for local EBM studies. This is justified because the computational patch remains small compared with the heliocentric distance, allowing the EBM mapping of local spherical expansion onto a Cartesian, periodic domain. Periodicity reduces computational cost and therefore enables higher numerical resolution to capture the turbulent cascade across a broad range of scales. We stress the limitations: periodic domains omit open-boundary effects (global inflow/outflow) and can box-limit the longest parallel modes (affecting $\ell_{\parallel}$ and near-2D structures). To mitigate these effects we use an initially anisotropic (elongated) domain so transverse stretching is naturally represented as $L_{y},L_{z}\propto a$. Within these bounds the model effectively captures the local, expansion-driven dynamics of reflection-driven turbulence beyond the Alfvén critical point.
We set the initial box dimensions to $L_{x0}=1$ and $L_{y0}=L_{z0}=0.2$. We use the expansion factor $a$ to parameterize the evolution of the system, where $a$ is directly related to heliocentric distance via $a=R/R_{0}$, with $R_{0}$ being the radial distance at which the simulation is initialized ($a=1$). The choice of $R_{0}$ is arbitrary in the sense that it sets the reference scale and simply rescales other quantities (e.g., $v_{\rm A}$) without changing the dimensionless physics. To aid interpretation and correspondence with observational regimes, we can map $a$ to actual solar distances using the evolving Parker spiral angle $\Phi(a)$. Assuming a Parker spiral angle of $45^{\circ}$ occurs at 1 AU (Borovsky, 2010). In our simulations with initial Parker spiral angle $\Phi_{0}=2^{\circ}$, the value $a\,\approx\,28.64$ corresponds to 1 AU, while for $\Phi_{0}=5^{\circ}$, we have $a\approx 11.5$ at 1 AU. Importantly, the mapping between the heliocentric distance and PS is not unique: it depends on the PS calibration (e.g. the reference angle at 1 AU), the latitudes, and on how one chooses the simulation inner boundary. At high latitudes, the PS is much weaker, so off-ecliptic regions remain nearly radial even out to several AU. Note that, scanning over $\Phi_{0}$ at fixed initial conditions is best interpreted as starting from the same initial $\chi_{\rm exp,0}$ (discussed below) at different heliocentric distances.
As the system expands, the box elongates in the transverse directions; by $a=50$, the domain becomes strongly oblate, resembling a pancake-like structure. At $a=1$, we initialize the background magnetic field $\bar{\bm{B}}$} in the $xy$ plane with $|{\bar{\bm{B}}}|=1$ and $B_{y,0}<0$, defining the initial Parker‐spiral angle $\Phi_{0}$.
@@
which measures how the components of $\bm{B}$ are correlated to keep $|\bm{B}|$ constant. $C_{B}^{2}=0$ for perfectly spherically polarized fluctuations. The initial condition is chosen to mimic a strongly $z^{+}$-dominated, nearly spherically polarized Alfvénic wind propagating outwards from near the Alfvén point: $\bm{z}^{+}(t=0)$ carries the fluctuation energy while $\bm{z}^{-}(t=0)=0$. Radial expansion and large-scale inhomogeneity produce partial reflection (creation of $\bm{z}^{-}$) as the packet propagates outward, thereby initiating reflection-driven turbulence. We initialize the fluctuations as linearly polarized Alfvén waves ($\delta\bm{u}=\delta\bm{B}/\sqrt{4\pi\rho}$, corresponding to $\bm{z}^{-}=0$) with $\delta\bm{u}$ and $\delta\bm{B}$ lying in the $(\bm{k}\times{{\bar{\bm{B}}}})$ direction (Squire et al., 2020; Johnston et al., 2022). The initial set of Alfvén waves are generated from a sum of randomly phased waves with amplitude determined by Gaussian energy spectrum,
@@
Here $k_{\parallel,0}=\kappa_{\parallel}({2\pi}/{L_{x}}),\;k_{\perp,0}=\kappa_{\perp}({2\pi}/{L_{\perp}}),\;\text{and}\;k_{w}=({12}/{L_{\perp}})$, which set the scale of the center of Gaussian peak and the width of the peak respectively. $k_{\parallel,0}$ and $k_{\perp,0}$ lie respectively parallel and perpendicular to $\bar{\bm{B}}$, maintaining anisotropy even with an initial tilted spiral field. As the expanding-box simulation begins, the modes rapidly adjust within approximately 0.5 Alfvén times, losing the variation in $|\bm{B}|$ caused by their initial linear polarization. This natural relaxation reduces the magnetic compressibility $C_{B}^{2}$ and allows the fluctuations to become spherically polarized.
To choose simulation parameters and quantify regimes, we use the timescales and their ratios, as discussed in section˜2. The expansion time is $\tau_{\rm exp}^{-1}=\dot{a}/a$, the linear timescale is $\tau_{\rm A}^{-1}=k_{\parallel}v_{\rm A}$, and the outer-scale nonlinear time is $\tau_{\rm nl}^{-1}=k_{\perp}z^{+}$. From these rates we form the dimensionless control parameters: (i) $\chi_{\rm exp,0}=(k_{\perp,0}\;z^{+}_{\rm rms0})/(\dot{a}/a)$; (ii) $\chi_{\rm A}\doteq(k_{\perp}z^{+})/(k_{\parallel}v_{\rm A})$, which compares strength of nonlinear interactions to linear Alfvénic propagation and thus quantifies how strongly nonlinearity competes with Alfvénic de-correlation of $z^{-}$ fluctuations (Goldreich and Sridhar 1995); and (iii) the ratio of box-scale Alfvén time to the expansion time $\Delta_{0}\doteq\chi_{\rm exp0}/\chi_{\rm A0}=v_{\rm A,0}/\dot{a}L_{x}$ which is constant. For the parameter set used in this paper we fix $\Delta_{0}=2$ (via $\dot{a}=0.5$, $L_{\parallel}=1$ and $v_{\rm A}=1$) and vary $\chi_{\rm exp,0}$ roughly in the range $\approx 11\!-\!30$. We explore two initial fluctuation amplitudes defined by $A\equiv{z^{+}_{\rm rms0}}/{v_{\rm A,0}}=1.0\;\text{and}\;0.5,$ where $z^{+}_{\rm rms0}$ is the initial root-mean-square fluctuation amplitude (see Table 1 for exact values). The moderate-resolution simulations with reduced amplitude ($A_{0}=0.5$) are designated as A05 to distinguish them from the standard moderate-resolution runs with $A_{0}=1.0$. Changing $A$ alters the initial nonlinear time (and therefore $\chi_{\rm exp,0}$ and $\chi_{\rm A,0}$), so we fix $\chi_{\rm exp,0}$ to isolate pure amplitude effects. Physically, larger $A$ produces larger fluctuations with larger $\delta B_{\parallel}$ to maintain their spherical polarization, whereas smaller $A$ (initial conditions) remain closer to the reduced MHD limit, which is derived assuming $z^{+}/v_{\rm A}\ll 1$.
@@ 4.1 Turbulence Evolution and Emergence of the Parker Spiral
The morphology of turbulence depends sensitively on the background magnetic geometry. The top panels of Fig. 3a depict two-dimensional snapshots ($y$-$z$ mid plane slice) of the perpendicular normalized Elsässer fields $|\bm{z}_{\perp}^{\pm}|$ for the radial A05-$\Phi_{0}\!=\!0^{\circ}$ run (here $\bm{z}_{\perp}^{\pm}=(\bm{I}-\hat{\bm{b}}\hat{\bm{b}})\cdot\bm{z}^{\pm}$, which is approximately the Alfvénic part). In the radial case, turbulence evolves in two distinct phases. During the early, imbalanced stage when $\chi_{\rm exp}>1$, outward $z^{+}$-fluctuations are strong and develop fine perpendicular structure, while reflected $z^{-}$ fluctuations appear as more intermittent, field-aligned filaments. As expansion proceeds and $\chi_{\rm exp}$ approaches unity, reflected normalized wave-action energy grows and $\tilde{E}^{-}/\tilde{E}^{+}$ approaches unity, with both fields acquiring increasingly similar circular vortex-like structures. In this balanced phase, the nonlinear transfer rate decreases relative to expansion and the cascade stalls, leaving coherent vortices to dominate the dynamics and shutting off heating. Throughout both stages, expansion stretches eddies across the mean field, generating large-aspect-ratio quasi 2D eddies towards the simulation’s end. This behavior qualitatively matches the reduced-MHD results of Meyrand et al. (2025), who reported 2D “Alfvén vortices”, similar to those we see, in the late stage $\chi_{\rm exp}\sim 1$ magnetically dominated regime.
@@
The Parker spiral simulations evolve differently. In A05-$\Phi_{0}\!=\!1.5^{\circ}$ (lower two panels of Fig. 3), turbulence does not relax toward a quasi-two-dimensional state as in the radial case. Instead, nonlinear activity persists to larger expansion factors because (as we will show below) $\chi_{\rm exp}$ remains larger. The fields develop persistent, ribbon-like structures with sharper transverse gradients and smaller $\ell_{\perp}$ than in radial runs (this will be quantified in section˜4.3). Physically, this follows because the continuous rotation of the mean field during expansion maintains a persistent misalignment between the perpendicular and expansion directions; as illustrated in Fig. 1 and discussed in section˜2.4.1, that misalignment prevents eddies from maintaining cylindrical symmetry about the local field and thereby inhibits the formation of quasi-2D vortices. Consequently, we observe a persistent, irregular 3D anisotropy in PS runs and persistent turbulent heating.
@@
Figure 4 shows similar features via a three-dimensional depiction of the Elsässer fields from the high-resolution (HR-$\Phi_{0}\!=\!0^{\circ}$ and HR-$\Phi_{0}\!=\!4^{\circ}$) simulations. These volumetric views support the conclusion of Fig. 3, and also show the evolution of radial structure. In HR-$\Phi_{0}\!=\!0^{\circ}$, turbulence condenses into slowly developing Alfvén vortices, though these are not so obvious here because of the earlier time. In HR-$\Phi_{0}\!=\!4^{\circ}$, structures stay at smaller scale and are possibly more intermittent, developing 3D outer-scale anisotropy. The resulting sustained small $\ell_{\perp}$ inhibits the cascade from stalling. We also see qualitative agreement with prior studies linking the PS geometry to 3D anisotropy (Verdini et al., 2019).